# Adversarial Disentanglement by Backpropagation with Physics-Informed Variational Autoencoder

### v1

@@ 5.3 Two-span bridge
The final synthetic case study utilizes the two-span bridge benchmark presented in [Tatsis2019], illustrated in Figure 1(c). Members of a homogeneous population [Bull2021] of bridges are subjected to controlled loading tests, where a vehicle with known mass and moving at a constant velocity is used to excite the bridge response. The response is obtained as a strain influence line, measured by a point strain gauge placed under the center of the left midspan of the bridge. Each time-series is composed of $64$ measurements, equally spaced in time $t\in[0,25]$ s. The behavior of each bridge is partially determined by the unknown vertical stiffnesses of the supports $k_{\mathrm{v},1}$, $k_{\mathrm{v},2}$ and $k_{\mathrm{v},3}$, which are taken to vary between different bridges due to variability in the design, construction and soil conditions. The boundary conditions are known to be symmetric such that $k_{\mathrm{v,1}}=k_{\mathrm{v,3}}$. The vertical stiffnesses are considered as physics-grounded latent variables. It is assumed that the position of the central pier can vary between members of the population by up to $2.0$ m. The nominal physics-based model includes the location of the pier as a parameter $\delta_{\mathrm{s}}$, describing the offset relative to the center of the bridge in the longitudinal direction. Furthermore, fluctuations $\delta_{\mathrm{v}}$ from the prescribed reference vehicle velocity $v_{\mathrm{ref}}$ were observed during the tests that can not be accounted for in the nominal physics-based model. Noisy measurements of the vehicle velocity, and the known pier offsets $\delta_{\mathrm{s}}$ are included as domain variables. During inspections performed by experts, the left, middle and right support of each bridge are assigned scores $\bm{y}=(y_{1},y_{2},y_{3})\in[0,1]^{3}$, with zero representing pristine condition and unity representing complete loss of the cross-section. These scores are considered as class observables. Finally, a small variability $\delta_{\mathrm{F}}$ in the reference vehicle load $F$ stemming from unknown variability in the transverse position of the vehicle is considered as an unknown confounding influence. The response, domain and class observables are contaminated with i.i.d. samples of Gaussian white noise with standard deviations $\sigma_{\mathrm{x}}=\sigma_{\mathrm{c}}=\sigma_{\mathrm{y}}=0.001$. The quantities involved in the case study along with their prior and ground truth distributions are summarized in Table 3. It is important to note that this case study is not intended to be a realistic procedure for system identification and SHM of bridges, and circumvents several important practical difficulties such as ensuring the consistency and alignment of data collected over long time-scales from a large number of structures. Furthermore, it is assumed for simplicity that the degradation condition of the bridges does not change significantly within the amount of time required to obtain the dataset.
@@
Variable Unit Type Prior distribution Ground truth
$\log k_{\mathrm{v},1}$ Pa Physical $\mathcal{U}(8.0,12.0)$ $\mathcal{U}(9.0,11.0)$
$\log k_{\mathrm{v},2}$ Pa Physical $\mathcal{U}(8.0,12.0)$ $\mathcal{U}(9.0,11.0)$
$y_{1}$ Class - $\mathcal{U}(0.0,0.9)$
$y_{2}$ Class - $\mathcal{U}(0.0,0.9)$
$y_{3}$ Class - $\mathcal{U}(0.0,0.9)$
$\delta_{\mathrm{v}}$ m/s Domain - $\mathcal{U}(-0.1,0.1)$
$\delta_{\mathrm{s}}$ m Domain - $\mathcal{U}(-2.0,2.0)$
$\delta_{\mathrm{F}}$ kN Unknown - $\mathcal{U}(-0.1,0.1)$
Training data for the full and nominal physical models is generated using the FE model of [Tatsis2019] as a simulator. The FE model is composed of quadrilateral isoparametric plane stress elements with $9$ nodes each, arranged in a $200\times 6$ grid. The length, width and height are assumed constant and equal to $L=12.5$ m, $w=0.1$ m and $h=0.6$ m for all of the bridges. The material is linear elastic, with Young’s modulus $E=210$ GPa, density $\rho=7800$ kg/m3 and Poisson’s ratio $\nu=0.3$. All supports are modeled as linear springs in the vertical direction and fully fixed in the horizontal direction. The equations of motion are integrated from $t_{0}=0$ seconds to $t_{1}=25$ seconds with a timestep of $\mathrm{d}t=0.00025$ seconds, using an implicit Newmark scheme with parameters $\gamma=1/2$ and $\beta=1/6$. It is assumed that the bridges are subject to deterioration in the form of a percent reduction of the cross section thickness, ranging from $0\%$ to $90\%$ in the region spanning $0.1\cdot L$ around each support. It is noted that the deterioration is intentionally exaggerated to an unrealistic extent in order to ensure that the damage can be observed in the strain influence lines. The damaged regions are illustrated in Figure 1(c). The corresponding nominal physics-based model is a simplistic but representative model for the behavior of the bridges in the population under study in their nominal condition. Given the vertical stiffness of the abutments and support, and the offset of the central pier, the nominal model returns a time-series of strains.
The model is trained with $\lambda=1/1024$ and $d_{z_{\mathrm{c}}}=d_{z_{\mathrm{y}}}=4$. The predictions generated by the model while traversing each of the generative factors are shown in Figure 10. In contrast to the damped oscillator case study (8(b)), here the influence of the unknown confounded $\delta_{\mathrm{F}}$ can be captured by the physics-based model. Furthermore, it can be seen that the adversarial training effectively constrains the decoder such that the data-driven component of the reconstructed response is constant when the domain and class generative factors are constant.
@@
Figure 10: Mean prediction and ±$\pm 2\sigma$ uncertainty bounds for the physics-based (top row) and data-driven (middle row) components, and combined prediction (bottom row) while traversing the generative factors. The input response measurements are denoted as dots in the bottom row.
The corresponding traversal of the latent space, shown in Figure 11, reveals that the variability in $\bm{x}$ is primarily captured by the physics-grounded subset of the latent space. The domain and class subsets of the latent variables encode information that enables the auxiliary decoders to predict the domain and class labels, and the response decoder to correct the physics-based model prediction. It can be seen that the influence of the unknown confounder $\delta_{\mathrm{F}}$ is captured as variability in the physics-based subset $\bm{z}_{\mathrm{x}}$, indicating that model form uncertainty is compensated by inferring an "effective" value of the physics-grounded latent variables. This is in contrast to the damped oscillator example (Section 5.2) where the influence of the unknown confounder could not be captured by the physics-based model, resulting in increased uncertainty in the decoder.
@@
As discussed in Section 3, the model is trained in a fully supervised manner to simultaneously reconstruct the domain and class variables from the input measurements, making it possible to handle tasks such as damage detection, where predicting the class labels $\bm{y}$ from input measurements $\bm{x}$ is of interest. Given a trained model, the condition labels of a similar bridge can be predicted from response measurements. The performance is evaluated in two different cases, illustrated in Figure 12, referred to as "interpolation" and "extrapolation" respectively. For each case, the space of physics-based generative factors is subdivided into four quarters. In the interpolation case, the model is trained on data from three quarters and evaluated on test data sampled from the fourth. In the extrapolation case, the model is trained on data from a single quarter and evaluated on the remaining three. To obtain a more comprehensive and representative evaluation, each of the two cases is divided into four sub-cases, over which the results are averaged.
@@
The proposed model using two different hyperparameter settings, denoted as DPIVAE-A and DPIVAE-B, is compared with linear regression (LIN), Gaussian Process Regression (GPR) and a Multi-Layer Perceptron (MLP). For DPIVAE-A the GRL is not utilized, i.e. $\lambda=-1$, and separate encoders are used for each subset of the latent variables. For DPIVAE-B the GRL hyperparameter is taken as $\lambda=1/1024$. The GPR is implemented with a Radial Basis Function kernel and additive Gaussian white noise. The MLP is formulated with two hidden layers, each with a width of $128$ units and a Rectified Linear Unit (ReLU) activation function. It is important to note that the results provided below are not a rigorous comparison of the proposed approach against existing models. Rather, they are intended to highlight that the performance is comparable, and that the model can feasibly be used to predict the class variables in a complex high-dimensional case study. Only limited manual hyperparameter tuning is performed for all the models involved in the comparison. The performance of each model is quantified in terms of the R2 and Mean Squared Error (MSE), shown in Table 4.
@@
DPIVAE-A 0.854 $\pm$ 0.027 0.010 $\pm$ 0.002 0.808 $\pm$ 0.038 0.013 $\pm$ 0.003
DPIVAE-B 0.904 $\pm$ 0.013 0.006 $\pm$ 0.001 0.866 $\pm$ 0.029 0.009 $\pm$ 0.002
GPR 0.922 $\pm$ 0.007 0.005 $\pm$ 0.000 0.889 $\pm$ 0.027 0.007 $\pm$ 0.002
LIN 0.766 $\pm$ 0.019 0.016 $\pm$ 0.001 0.724 $\pm$ 0.020 0.018 $\pm$ 0.001
MLP 0.787 $\pm$ 0.027 0.014 $\pm$ 0.002 0.735 $\pm$ 0.050 0.018 $\pm$ 0.003
The proposed approach using adversarial training doesn’t result in a noticeable improvement over existing approaches for the specific interpolation and extrapolation tasks in the present case study. The GPR outperforms all other models, including the DPIVAE. This result is particularly unexpected considering that it only uses a fraction of the parameters. In both cases the model performs better when adversarial training is used. The analysis indicates that the proposed approach performs on par with commonly used data-driven approaches, but with the added benefit of ensuring the proper use of the known physics and the additional interpretability of the physics-grounded latent space. Importantly, Figure 10 and Figure 11 illustrate that the adversarial training can feasibly constrain $g_{\bm{\theta}}(\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$, such that it only contributes to the prediction when justified by additional domain and class observables. We note that the proposed approach outperforms all other models on the same interpolation and extrapolation tasks in the Euler-Bernoulli beam example. The investigation of the factors that affect the performance of the proposed approach, and the conditions under which the incorporation of the known physics and the adversarial training can provide a benefit in class prediction tasks as an avenue for future work.
@@ 6 Discussion
The results presented in Section 5 indicate that the proposed architecture and adversarial objective effectively constrain the posterior distribution over domain and class latent variables, and by extension, the flexibility of the data-driven decoder components. The strength of the regularization is controlled by an interpretable and intuitive hyperparameter that determines the strength of the gradient reversal. This allows for the main and auxiliary decoders to be trained in a collaborative or adversarial manner. This hyperparameter effectively controls the relative importance of the physics-based and data-driven components, and can be used to encourage the model to preferentially utilize the known physics. When the training is adversarial, the domain and class latent spaces encode features of the response measurements that can be related to the domain or class variables, and that can not be accounted for by the known physics, and are simultaneously constrained to avoid overriding the physics-based model predictions. Because neither the domain or class observables are necessary during model evaluation, the proposed approach has the potential to reduce the need for cumbersome and expensive data collection methods, such as those involving elaborate experimental procedures or expert assessments.
@@
It is important to consider the assumptions and limitations of the proposed approach. One of the main drawbacks of the model is the additivity assumption imposed on the physical, domain and class components of the response. It is expected that the model will perform sub-optimally when this assumption is violated. Furthermore, the accuracy of the inferred physics-grounded latent variables will depend on the relative contribution of the physics, domain and class influences to the measured response. Significant domain and class contributions to the response, or violating the additivity assumption of Equation 2.3, can lead to inaccurate inference of physics-grounded latent variables and large uncertainty in the predictions. Additionally, the model requires that multiple types of data are available. This includes measurements of the structural response, and information on domain and class. In SHM applications, this might necessitate data alignment procedures of response measurements, environmental conditions and damage level descriptions, and could potentially limit the immediate applicability of the proposed architecture. It is not possible to provide a rule-of-thumb about the amount of data required for effective training. This would be dependent on the specific problem, noise levels, physics-based model, and accuracy of the domain and class information, and also on the particular architecture choices (e.g. the number and depth of layers used in the feed-forward NNs in the encoder and decoder). In some applications, the incorporation of the known physics might lead to a reduction in the data requirements. This has not been investigated in the current paper, since it is believed that it would be strongly dependent on the case study chosen, rather than offering any general insight. It is worth mentioning that the interaction between the encoder, decoders, the GRL, and the known physics can be unintuitive in some applications, limiting the applicability of the approach and potentially necessitating implicit supervision by a human expert. Finally, it is known that adversarial training can result in numerical instability, e.g. in GANs [Wiatrak2020]. Throughout this work, occasional instability and oscillatory behavior were observed during training. We found that these issues could generally be addressed by adjusting the $\lambda$ hyperparameter, and by implementing early stopping based on the value of the ELBO on a held-out validation set.

### v2

@@ 5.3 Bridge case study
The final synthetic case study utilizes the two-span bridge benchmark presented in [Tatsis2019], illustrated in Figure 1(c). Members of a homogeneous population [Bull2021] of bridges are subjected to controlled loading tests, where a vehicle with known mass and moving at a constant velocity is used to excite the bridge response. The response is obtained as a strain influence line, expressed in parts per thousand (‰), measured by a point strain gauge placed at a distance of $5.625$ m from the start of the bridge, and at a height of $0.1$ m from the bottom of the cross section. Each time-series is composed of $64$ measurements, equally spaced in time $t\in[1,21]$ s, where $t=0$ s is the moment the vehicle enters the bridge. The behavior of each bridge is partially determined by the unknown vertical stiffnesses of the supports $k_{\mathrm{v},1}$, $k_{\mathrm{v},2}$ and $k_{\mathrm{v},3}$, which are taken to vary between different bridges due to variability in the design, construction and soil conditions. The boundary conditions are known to be symmetric such that $k_{\mathrm{v,1}}=k_{\mathrm{v,3}}$. The base-10 logarithms of the vertical stiffnesses are considered as physics-grounded latent variables. In the horizontal direction, only the left support has a large stiffness, while the rest are unconstrained. It is assumed that the position of the central pier can vary between members of the population by up to ±$\pm 1.0$ m from $L/2$. The nominal physics-based model includes a parameter $\delta_{\mathrm{s}}$, describing the offset of the pier relative to the center of the bridge in the longitudinal direction. Furthermore, fluctuations from the prescribed reference vehicle velocity $v_{\mathrm{ref}}=1$ m/s were observed during the tests that can not be accounted for in the nominal physics-based model. These fluctuations are modeled as a multiplicative term $\delta_{\mathrm{v}}$ such that $v=\delta_{\mathrm{v}}\cdot v_{\mathrm{ref}}$. Noisy measurements of the vehicle velocity, and the known pier offsets $\delta_{\mathrm{s}}$ are included as domain variables. It is assumed that the bridge decks are prone to deterioration in a region around the supports. During inspections performed by experts, each bridge is assigned scores $\bm{y}=(y_{1},y_{2})$, quantifying the deterioration of the structure near the left and middle supports respectively. Each $y_{i}$ takes values between zero and one, where zero represents pristine condition and unity corresponds to severe damage of the cross-section at that position. These scores are considered as class observables. Finally, a small variability $F_{\mathrm{ref}}$ is considered in the vehicle load such that $F=\delta_{\mathrm{F}}\cdot F_{\mathrm{ref}}$, where $F_{\mathrm{ref}}=100$ kN is the reference load. This variability is caused by deviation in the transverse position of the vehicle, and is considered as an unknown confounding influence. The response, domain and class observables are contaminated with i.i.d. samples of Gaussian white noise with standard deviations $\sigma_{\mathrm{x}}=\sigma_{\mathrm{c}}=\sigma_{\mathrm{y}}=10^{-4}$. The quantities involved in the case study along with their prior and ground truth distributions are summarized in Table 3. It is important to note that this case study is not intended to be a realistic representation of system identification and SHM for bridges, since it circumvents several important practical difficulties such as ensuring the consistency and alignment of data collected over long time-scales from a large number of structures. Furthermore, it is assumed for simplicity that the degradation condition of the bridges does not change significantly within the amount of time required to obtain the dataset.
@@
Variable Unit Type Prior distribution Ground truth Reference value
$\log_{10}k_{\mathrm{v},1}$ N/m Physical $\mathcal{U}(9.0,12.0)$ $\mathcal{U}(9.5,11.5)$ 11.5
$\log_{10}k_{\mathrm{v},2}$ N/m Physical $\mathcal{U}(9.0,12.0)$ $\mathcal{U}(9.5,11.5)$ 11.5
$y_{1}$ Class - $\mathcal{U}(0.0,1.0)$ 0.1
$y_{2}$ Class - $\mathcal{U}(0.0,1.0)$ 0.1
$\delta_{\mathrm{v}}$ Domain - $\mathcal{U}(0.9,1.1)$ 1.0
$\delta_{\mathrm{s}}$ m Domain - $\mathcal{U}(-1.0,1.0)$ 0.0
$\delta_{\mathrm{F}}$ Unknown - $\mathcal{U}(0.9,1.1)$ 1.0
Training data for the full and nominal physical models is generated using the FE model of [Tatsis2019] as a simulator. The FE model is composed of quadrilateral isoparametric plane stress elements with $9$ nodes each, arranged in a $200\times 6$ grid. The length, width and height are assumed constant and equal to $L=25.0$ m, $w=0.1$ m and $h=0.6$ m for all of the bridges. The material is linear elastic, with Young’s modulus $E=200$ GPa, density $\rho=7850$ kg/m3 and Poisson’s ratio $\nu=0.3$. All supports are modeled as linear springs in the vertical direction. The equations of motion are integrated from $t_{0}=0$ seconds to $t_{1}=25$ seconds with a timestep of $\mathrm{d}t=0.00025$ seconds, using an implicit Newmark scheme with parameters $\gamma=1/2$ and $\beta=1/6$. The deterioration is modeled as a reduction of the cross section width, ranging from $0\%$ (for $y_{i}=0.0$) to $90\%$ (for $y_{i}=1.0$) in the region spanning $L/20$ around the corresponding support. It is noted that the deterioration is intentionally exaggerated to a large extent to ensure that it can be observed in the strain influence lines. The damaged regions are illustrated in Figure 1(c). The nominal physics-based model is assumed to be a simplistic but representative model for the behavior of the bridges in the population under study in their nominal condition, obtained through the procedure described in Section 5. Given the vertical stiffness of the abutments and support, and the offset of the central pier, the nominal model returns a time-series of strains.
The model is trained with $\lambda=1/1024$ and $d_{z_{\mathrm{c}}}=d_{z_{\mathrm{y}}}=4$. The predictions generated by the model while traversing each of the generative factors are shown in Figure 10. It can be seen that the data-driven component of the decoder is prevented from capturing variability in the reconstructed response when varying $\log_{10}k_{\mathrm{v},1}$, $\log_{10}k_{\mathrm{v},2}$ and $\delta_{\mathrm{F}}$, but is able to contribute to the components caused by the variation of the domain and class generative factors. Furthermore, the figure illustrates that the unknown confounder $\delta_{\mathrm{F}}$ can be partially accounted for by the physics based model. This is in contrast to the oscillator example (Section 5.2) where the influence of the unknown confounder could not be accounted for by the known physics.
@@
Figure 10: Mean prediction and ±$\pm 2\sigma$ uncertainty bounds for the physics-based $\hat{\bm{x}}_{\mathrm{p}}$ and data-driven $\hat{\bm{x}}_{\mathrm{d}}$ components, and combined prediction $\hat{\bm{x}}$ while traversing the generative factors. The input response measurements are denoted as dots in the bottom row.
The previous conclusions are further supported by the traversal of the latent space, shown in Figure 11, which indicates that the domain and class subsets of the latent variables encode information that enables the auxiliary decoders to predict the domain and class labels, and the response decoder to correct the physics-based model prediction. It can be seen that the influence of the unknown confounder $\delta_{\mathrm{F}}$ is partly captured as variability in the physics-based subset $\bm{z}_{\mathrm{x}}$, indicating that model form uncertainty is compensated by inferring an “effective” value of the physics-grounded latent variables. Figure 11 also suggests that the domain and class latent variables only capture variability in the corresponding generative factors, whereas the physics-based latent variables are always active, providing additional evidence for the claim that the adversarial objective induces a disentangled representation while prioritizing the use of the known physics. Importantly, Figure 10 and Figure 11 illustrate that the adversarial training can feasibly constrain $g_{\bm{\theta}}(\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$, such that it only contributes to the prediction when justified by additional domain and class observables.
@@
As discussed in Section 3, the model is trained in a fully supervised manner to simultaneously reconstruct the domain and class variables from the input measurements, making it possible to handle tasks such as damage detection, where predicting the class labels $\bm{y}$ from input measurements $\bm{x}$ is of interest. Given a trained model, the condition labels of a similar bridge can be predicted from response measurements. The performance is evaluated in two different cases, illustrated in Figure 12, referred to as “interpolation” and “extrapolation” respectively. For each case, the space of physics-based generative factors is subdivided into four quarters. In the interpolation case, the model is trained on $N_{\mathrm{train}}=1024$ samples from three quarters and evaluated on $N_{\mathrm{test}}=512$ samples from the fourth. In the extrapolation case, the model is trained on data from a single quarter and evaluated on the remaining three, using the same train and test set sizes. All other generative factors are sampled from the ground truth distributions presented in Table 3. To obtain a more comprehensive evaluation, each of the two cases is divided into four sub-cases, over which the results are averaged.
@@
The proposed model using two different hyperparameter settings, denoted as DPIVAE-A and DPIVAE-B, is compared with linear regression (LIN), Gaussian Process Regression (GPR) and a Multi-Layer Perceptron (MLP). For DPIVAE-A the GRL is not utilized, i.e. $\lambda=-1$, and separate encoders are used for each subset of the latent variables. For DPIVAE-B the GRL hyperparameter is taken as $\lambda=1/1024$. The GPR is implemented with a Radial Basis Function kernel and additive Gaussian white noise. The MLP is formulated with two hidden layers, each with a width of $64$ units and a Rectified Linear Unit (ReLU) activation function. Results in terms of interpolation and extrapolation performance of each model is quantified in terms of the R2 and Mean Squared Error (MSE), shown in Table 4. These results are intended to highlight that the performance of the different models is comparable, and that the proposed model can feasibly be used to predict the class variables in a complex high-dimensional case study. Manual hyperparameter tuning is performed for the models involved in the comparison.
@@
DPIVAE-A 0.905 $\pm$ 0.023 0.008 $\pm$ 0.002 0.676 $\pm$ 0.153 0.027 $\pm$ 0.013
DPIVAE-B 0.943 $\pm$ 0.012 0.005 $\pm$ 0.001 0.809 $\pm$ 0.090 0.016 $\pm$ 0.008
GPR 0.957 $\pm$ 0.022 0.004 $\pm$ 0.002 0.820 $\pm$ 0.112 0.015 $\pm$ 0.009
LIN 0.863 $\pm$ 0.005 0.011 $\pm$ 0.000 0.617 $\pm$ 0.285 0.032 $\pm$ 0.024
MLP 0.839 $\pm$ 0.035 0.013 $\pm$ 0.003 0.365 $\pm$ 0.397 0.053 $\pm$ 0.033
The proposed approach using adversarial training doesn’t result in any improvement over existing approaches for the specific interpolation and extrapolation tasks in the present case study. This can be attributed to the adversarial training, which forces the encoder to weigh any information that is relevant to the prediction of $\bm{y}$ against the potential improvement it provides towards the reconstruction of $\bm{x}$. It can be seen that the GPR outperforms all other models while only using a fraction of the parameters, possibly due to the smoothness of the input influence line measurements. The results also indicate that the model performs better when using a single encoder combined with adversarial training in both interpolation and extrapolation. Despite this negative result, we speculate that the disentangled representation induced by the architecture, the invariance of the class latent variables to unknown confounding influences, and the incorporation of the known physics, might be beneficial in certain tasks. Future work will aim to investigate the factors that affect the performance of the proposed approach, and the conditions under which it can provide a benefit in class prediction tasks. This analysis indicates that the proposed model performs on par with other commonly used data-driven models, but with the added benefit of ensuring the proper use of the known physics and the additional interpretability of the physics-grounded latent space.
@@ 6 Discussion
The results presented in Section 5 indicate that the proposed architecture and adversarial objective effectively constrain the posterior distribution over domain and class latent variables, and by extension, the flexibility of the data-driven decoder components. The strength of the regularization is controlled by an interpretable hyperparameter that determines the strength of the gradient reversal. This allows for the main and auxiliary decoders to be trained in a collaborative or adversarial manner. This hyperparameter effectively controls the relative importance of the physics-based and data-driven components, and can be used to encourage the model to preferentially utilize the known physics. When the training is adversarial, the domain and class latent spaces encode features of the response measurements that can be related to the domain or class variables, and that can not be accounted for by the known physics, and are simultaneously constrained to avoid overriding the physics-based model predictions. Because neither the domain or class observables are necessary during model evaluation, the proposed approach has the potential to reduce the need for cumbersome and expensive data collection methods, such as those involving elaborate experimental procedures or expert assessments.
@@
It is important to consider the assumptions and limitations of the proposed approach. One of the main drawbacks of the model is the additivity assumption imposed on the physical, domain and class components of the response. It is expected that the model will perform sub-optimally when this assumption is violated. Furthermore, the accuracy of the inferred physics-grounded latent variables will depend on the relative contribution of the physics, domain and class influences to the measured response. Significant domain and class contributions to the response, or violating the additivity assumption of Equation 2.3, can lead to inaccurate inference of physics-grounded latent variables and large uncertainty in the predictions. Additionally, the model requires that multiple types of data are available, namely measurements of the structural response and information on domain and class. In SHM applications, this might necessitate data alignment procedures of response measurements, environmental conditions and damage level descriptions, and could potentially limit the immediate applicability of the proposed architecture. It is not possible to provide a rule-of-thumb about the amount of data required for effective training. This would be dependent on the specific problem, noise levels, physics-based model, and accuracy of the domain and class information, and also on the particular architecture choices (e.g. the number and depth of layers used in the feed-forward NNs in the encoder and decoder). In some applications, the incorporation of the known physics might lead to a reduction in the data requirements. This has not been investigated in the current paper, since it is believed that it would be strongly dependent on the case study chosen, rather than offering any general insight. It is worth mentioning that the interaction between the encoder, decoders, the GRL, and the known physics can be unintuitive in some applications, limiting the applicability of the approach and potentially necessitating implicit supervision by a human expert. Finally, it is known that adversarial training can be unstable [Wiatrak2020]. Throughout this work, occasional instability and overfitting were observed when using small datasets and large batch sizes. Depending on the case study and the value of the $\lambda$ hyperparameter, oscillatory behavior may also occur. We found that these issues could be addressed by adjusting the $\lambda$ hyperparameter, implementing early stopping based on the value of the ELBO on a held-out validation set, and reducing the batch size.